A courier service company has found that their delivery time of parcels to clients is approximately normally distributed with a mean delivery time of 30 minutes and a variance of 25 minutes (squared).

Required:

a) What is the probability that a randomly selected parcel will take longer than 33 minutes to deliver?

b) What is the probability that a randomly selected parcel will take less than 26 minutes to deliver?

c) What is the minimum delivery time (minutes) for the 2.5% of parcels with the longest time to

deliver?

d) What is the maximum delivery time (minutes) for the 10% of the parcels with the shortest time to deliver?


1
Expert's answer
2022-05-17T12:49:13-0400

Let "X=" delivery time: "X\sim N(\mu, \sigma^2)."

Given "\mu=30min, \sigma^2=25{min}^2"

a)


"P(X>33)=1-P(Z\le \dfrac{33-30}{\sqrt{25}})"

"=1-P(Z\le0.6)\approx0.2743"

b)


"P(X<26)=P(Z<\dfrac{26-30}{\sqrt{25}})"

"=P(Z<-0.8)\approx0.2119"


c)


"P(Z>\dfrac{x-30}{\sqrt{25}})=0.025"

"\dfrac{x-30}{5}=1.96"

"x=30+5(1.96)"

"x=39.8\ min"

d)


"P(Z<\dfrac{x-30}{\sqrt{25}})=0.10"

"\dfrac{x-30}{5}=-1.28155"

"x=30+5(-1.28155)"

"x=23.6\ min"